Reports 1-1 of 1 Clear search Modify search
fabian.arellano - 11:41 Sunday 21 February 2021 (16073) Print this report
Comment to Picomotor of PR3_BF_T(Motor2) got stuck (15976)

With Hirata-san, Washimi-san and Ikeda-san.

This is a report about the work carried out on Friday 19th of February.

  • We checked a couple of cables that looked touching each other in a picture, nevertheless, thay are not touching. Such an impression was created by the angle in which the picture was taken. See this picture and this other one.

  • We measured the resistance of H1 OSEM coil. It is 15.4 Ohm. I had the impresion there was a problem with the cable because the flag didn't seem to move when applying coil-magnet actuation.
  • We assesed the position of the OSEM flags. See this diagram.

Search Help